If an investor buys shares in a no-load mutual fund for $33 and after five years the shares appreciate to $57, what is (1) the percentage return and (2) the annual compound rate of return using time value of money? Round your answers to two decimal places. Percentage return: 0.7273 % The annual compound rate of return: %
Answer #1

The Percentage Return =(57-33)/33 =72.7273% or 72.73%
Annual compound rate of return =(1+72.7273%)^(1/5)-1 =11.55%